6 A monopolist has demand and cost curves given by: (5 marks) QD = 10,000 - 20P TC = 1,000 + 10Q + .05Q2 


1
Expert's answer
2021-07-16T09:36:39-0400

Given

"QD = 10,000 - 20P"

"TC = 1,000 + 10Q + .05Q^2"


"Q= 10,000 - 20P\\20P=10,000-Q\\P=\frac{10,000}{20}-\frac{Q}{20}"


"P=500-\frac{Q}{20}"


"TR=P\times Q\\"


"TR=(500-\frac{Q}{20})Q=500Q-\frac{Q^2}{20}"


"MR=500-\frac{Q}{10}"


"TC = 1,000 + 10Q + .05Q^2 \\MC=10+0.1Q"


"MR=MC\\500-\frac{Q}{10}=10+0.1Q\\500-10=0.1Q+\frac{Q}{10}\\490=\frac{Q}{5}\\Q=490\times5=2450"


"Q = 10,000 - 20P\\2450=10,000 - 20P\\20P=10,000-2450\\20P=7550\\P=377.5"